# Sky localization of gravitational waves from eccentric binaries

### v1

@@ Abstract
Orbital eccentricity in a compact binary inspiral is yet to be observed. We demonstrate that the orbital eccentricity in compact binary mergers can be used to improve their sky localization using gravitational wave observations. Existing algorithms that conduct the localizations are not optimized for eccentric sources. We use a semi-Bayesian technique to carry out localizations of simulated sources recovered using a matched-filter search. Through these simulations, we find that if a non-negligible eccentricity is obtained during the detection, an eccentricity-optimized algorithm can significantly improve the localization areas compared to the existing methods. The potential impact on the early-warning localization is investigated. We lay out the foundation for an eccentric early-warning system using the matched-filter search. We find cases of striking improvements while accounting for eccentricity toward potential eccentric neutron star binaries in the forthcoming observing scenarios of ground-based detectors. Improved localizations can be useful in effectually utilizing the capabilities of the follow-up facilities.
@@ I Introduction
Although a non-neglibible orbital eccentricity associated with a binary neutron star inspiral is expected to be a rare case, there are considerable efforts toward detecting any such systems (Huerta and Brown, 2013; Nitz et al., 2020a; Pal and Nayak, 2024; Dhurkunde and Nitz, 2023). As the low-frequency sensitivities of the current gravitational wave detectors improve, the chances of observing any eccentric binary inspiral become prominent and promising.
@@
Since orbital eccentricity has a prominent effect primarily in the low signal frequencies, we investigate its effects on the localization during the detection of a binary inspiral. Through simulations, we set up a scenario where mock signals from eccentric BNSs are recovered from simulated data using an eccentric matched-filter search and are then localized using an eccentricity-optimized localiztion algorithm. We primarily compare the outcome of the eccentric analysis with an, otherwise same, non-eccentric analysis. We show that in the presence of a non-negliglible eccentricity in the signals, the localizations can improve significantly. The improvement is expected to be more pronounced for lighter binaries where the number of low-frequency inspiral cycles in the sensitive frequency band is larger. In the absence of a residual eccentricity in the signals, the accuracy of the method falls back to that of the non-eccentric localizations on average.
The basic principle of the localization methods used in this work is to assume that the intrinsic source parameters obtained from matched-filter searches are the true values themselves  [^1] Alternately, this assumption can be lifted toward a full Bayesian parameter estimation, thus increasing the parameter dimensionality and the cost of computing the likelihoods. Several optimized approaches already implement this idea without including eccentricity (Morisaki and Raymond, 2020; You et al., 2021; Dax et al., 2021; Pathak et al., 2024).. In practice, the detection algorithms provide point estimates of these parameters that can carry large uncertainties, e.g. for eccentric sources (Pal and Nayak, 2024). Ideally, the set of true source parameters maximizes the recovered signal-to-noise ratio (SNR), however, as long as the any other combination forming a detection candidate, which nearly maximizes the SNR, can still be used. It has been shown that such a combination can be utilized for the purpose of rapid localization, laying the basis for the BAYESTAR algorithm which produces skymaps as accurate as that of a full parameter estimation. This has been demonstrated for sources with small aligned-spins (Singer and Price, 2016).
@@ II Simulated searches
This section briefly describes the processes used to recover simulated signals from simulated data. We create a simulated population of BNS systems with the signal parameters given in Table 1. To generate the signals, we use the TaylorF2Ecc waveform model (Moore et al., 2016). For any given detector, a signal is projected from the random sky location and is deposited in Gaussian noise colored with the simulated PSD for the detector. To recover injections in this work, we broadly make the following two broad changes to the PSO-based detection framework described in (Pal and Nayak, 2024). These changes are made primarily to simplify simulating the search process and to efficiently tackle a larger search parameter space.
@@
Figure 2: The coincident SNRs recovered by the PSO searches for simulated signals in the HL network having O4-like sensitivities. The searches are carried out in the frequency range of 20-512 Hz using the same datasets containing signals from spin-aligned eccentric BNSs immersed in Gaussian noise. We use a fixed template-sampling for both the searches as described in the text. The eccentric search recovers all the injections that are also recovered by the aligned-spin search. The commonly found injections are shown here.
Figure 3: Accuracy (left) and precision (right) of the localization areas compared for the simulated eccentric sources. The solid curves are obtained using the estimated intrinsic source parameters with the PSO-based searches. The dashed curves are obtained using the injected source parameters. The probability plots on the right show that the credible sky areas are consistent at a 99.7% confidence.
@@
Figure 4: Skymaps for an eccentric injection obtained with the aligned-spins only (left) and the eccentricity-optimized (right) localization algorithms. The injection is a randomly chosen (1.93, 1.88) $M_{\odot}$ source located at about 65 Mpc, having nomimal aligned-spins and an eccentricity of 0.1. The localizations are computed following sections Sec. II and III with simulated HLV network. The true location of the simulated source is maked with a star. Notably, we observe that the localization areas at the given credible levels are significantly improved with the eccentricity-optimized algorithm. Further, note the existence of two probability modes (left) compared to a single mode (right) at 90% credible level. Tracking skymaps with multiple separated modes could be resource-demanding process in the EM follow-up communities.
@@ III Localization algorithm
Once a detection has been made, the strain data around the event time from the observing detectors become of particular interest. Here, the localization algorithm receives a sufficient duration of the strain data containing the entire in-band signal determined by its chirp time. The chirp time is calculated using the chirp mass estimate of the source during its detection and the low-frequency cutoff. Note that the PSO search optimizes the SNR computed for the detector network for detection while including eccentricity as a search parameter. As indicated in (Singer and Price, 2016), the localization uncertainty is only semi-dependent on the uncertainties of the intrinsic source parameters. Thus we conduct a semi-Bayesian analysis by fixing the intrinsic parameters to that obtained as point estimates from the optimized template. Given the above information, it is possible to assign a probability that the source is located at any given part of the sky.
After fixing all the intrinsic parameters, we are left with the extrinsic parameters as shown in the Table. 1. We stochastically draw posterior samples from the following set of parameters,
@@
Note that the parameters given in Eq. 1 are intended to aid electromagnetic follow-up efforts. We use a uniform prior for each of the parameters. We choose to marginalize the remaining extrinsic parameters. Here it is performed analytically for $\varphi$ and numerically for $t_{c}$ and $\psi$ with a uniform grid of $1000$ points in the prior space. The posterior samples are obtained with dynesty sampler configured with nlive $=1000$ and dlogZ $=0.1$. This is accomplished within $25$ seconds using $\sim 32$ CPU cores. Note that these samples are obtained in the flat $\alpha-\delta$ coordinates. To compare with the localizations from BAYESTAR (Singer and Price, 2016), the samples are rendered into multi-ordered HEALPix format (Fernique et al., 2015; Górski et al., 2005; Zonca et al., 2019). Assigning probabilities to the pixels tiling the sky using the distribution of samples requires an additional postprocessing step  [^4] https://lscsoft.docs.ligo.org/ligo.skymap/tool/ligo_skymap_from_samples.html.
@@
Figure 6: Illustration of early-warning skymaps for an eccentric injection. Here the injected eccentricity is 0.05 for a (2.76, 1.43) $M_{\odot}$ source represented with a star. The left and the right panels show the aligned-spin and the eccentricity-informed skymaps; top, middle and bottom panels show the skymaps obtained till 32, 64 and 512 Hz respectively. Given the initial probability densities of the aligned-spin skymaps (the most probable locations are shown with solid pluses), a telescope should begin tracking at the lower left part in the skymap. Eventually, the telescope needs to be completely rotated as the final localization points toward a significantly different location, with little time left for the merger. The eccentric localizations do not require significant repointings in this case and the most probable locations are relatively more accurate. Here the quoted confidence areas are largely similar in both types of skymaps but we note the existence of such an effect in addition to the cases where the localization areas are themselves numerically improved. A similar effect is reported for non-eccentric sources in (Tohuvavohu et al., 2024).
[…]
To assess the precision of the localizations, we use a probability-probability (pp) plot to describe the credibility of the localization areas on the skymaps. A given credible localization patch on the skymap is determined by the group of pixels whose probabilities add up to the desired credible level, starting at the most probable pixel. The pp-plot is calculated by obtaining the fractions of the localized population containing the true location at all credible levels. We note that each configuration of localization described above requires a correction factor on the likelihood computations to obtain a self-consistent pp-plot discussed further in Appendix A. Thus, on average, the localization areas obtained with these methods can be considered reliable for the simulated population.
@@
Table 2: Approximate runtimes for generating event candidates from 256 seconds of Gaussian HL data stretch using the spin-aligned eccentric PSO search. The runtimes for CPU and GPU are obtained on a single AMD EPYC 9554 and NVIDIA Tesla V100 respectively utilizing all available cores. The GPU implementation is currently optimized for early-warning analyses only. The implementation of the early-warning localizations is not optimized and thus, have similar runtimes as for the full-bandwidth localizations.
Figure 7: Difference between the injected $t_{c}$ and the mean of the times-of-arrival of the simulated signals at H1 and L1. The number of injections are out of the 10000 injections when the SNR of the triggers crossed 7 at both the detectors. It is highly unlikely that Gaussian noise can surpass this arbitrary threshold in any detector. The plot indicates that a coincident window of 15 ms is insufficient to capture all coincidences in the detector triggers of astrophysical origin.
@@ IV Early-warning scenarios
Figure 8: Variation of the loss of recovered SNR by a simulated aligned-spin search (10-20 Hz) with the injected eccentricity and the total mass of the system in O5-like sensitivity.
Figure 9: Average of the early-warning SNRs lost by various searches for the commonly recovered injections with O5-projected HL sensitivities as a function of the early-frequency bands. We have also included the results for an otherwise same, eccentric-only EW search.
Figure 10: Accuracy (left) and precision (right) of the localization areas compared for the simulated aligned-spin sources. We note that in the absence of eccentricity in the signals, the localizations accuracy of the current technique is comparable to that of BAYESTAR. Similar to Fig. 3, the localization uncertainties are shown to be fairly self-consistent at 99.7% confidence.
@@
The commonly found injections are then localized in a similar fashion as described in Section III. The accuracies of the localization areas are compared for the observing scenarios as shown in Fig. 5. We observe that the eccentric localization is consistently more accurate than the non-eccentric one in all cases. However, the improvements are generally more prominent in O5-like sensitivities. We primarily attribute this to the greater SNR gain from the numerous low-frequency inspiral cycles while considering eccentricity at the detection stage. We argue that at very low data sampling rates, the SNR consistency across the detectors would carry relatively significant localization information, given the large errors in time and related phase measurements. An interesting comparison of the skymaps at different sampling frequencies is illustrated in Fig. 6.
In the above discussion, we have assumed a 15 ms time-window to test coincidences between triggers from LIGO Hanford (H1) and LIGO Livingston (L1). However, since the strain data for the EW analyses are critically downsampled, the timing uncertainty of the triggers of even astrophysical origin can result into non-coincidence, though they pass an SNR theshold criterion. This is demonstrated for the eccentric injections recovered with SNRs of at least 7 in both detectors as shown in Fig. 7. Refer to Appendix B for investigations on the background event rates resulting from increased coincidence window. We further observe that while an eccentric recovery obtains unbiased timing measurements on average, the aligned-spin only searches can report triggers with larger timing errors and/or with bias. Such errors are expected to propagate into the localizations which we do not explore here further.

### v4

@@ Abstract
We demonstrate that the orbital eccentricity in compact binary mergers can be used to improve their sky localization using gravitational wave observations. Existing algorithms that conduct the localizations are not optimized for eccentric sources. We use a semi-Bayesian technique to carry out localizations of simulated sources recovered using a matched-filter search. Through these simulations, we find that if a non-negligible eccentricity is obtained during the detection, an eccentricity-optimized algorithm can significantly improve the localization areas compared to the existing methods. We also lay out the foundation for an eccentric early-warning system using the matched-filter search. The potential impact on the early-warning localization is investigated. We indicate a few possible cases of improvements while accounting for eccentricity toward any detectable eccentric neutron star binaries in the forthcoming observing scenarios of ground-based detectors. Improved localizations can be useful in effectually utilizing the capabilities of the follow-up facilities.
@@ I Introduction
Although a non-neglibible orbital eccentricity associated with a binary neutron star inspiral is expected to be a rare case, there are considerable efforts toward detecting any such systems (Huerta and Brown, 2013; Nitz et al., 2020a; Pal and Nayak, 2024; Dhurkunde and Nitz, 2023). There are also several attempts to measure the eccentricities of LIGO-Virgo-KAGRA events, for example (Lenon et al., 2020; Romero-Shaw et al., 2020a; Gayathri et al., 2022; Gupte et al., 2024; Wu et al., 2020; Romero-Shaw et al., 2019, 2021, 2022, 2020b; Bonino et al., 2023; Gamba et al., 2023; Iglesias et al., 2022). As the low-frequency sensitivities of the detectors improve, the chances of observing any eccentric binary inspiral become prominent and promising, especially with the next-generation gravitational wave detectors (Dhurkunde and Nitz, 2023; Nitz and Wang, 2021). Studies indicating various observational prospects with eccentric binaries can be found in the literature, for example in (Yang et al., 2024a, 2022; Ma et al., 2017; Yang et al., 2024b).
@@
Since orbital eccentricity has a prominent effect primarily in the low signal frequencies, here we investigate its effects on the localization during the detection of a binary inspiral. Through simulations, we set up a scenario where mock signals from eccentric BNSs are recovered from simulated data using an eccentric matched-filter search and are then localized using an eccentricity-optimized localization algorithm. We primarily compare the outcome of the eccentric analysis with an, otherwise same, non-eccentric analysis. We show that in the presence of a non-negliglible eccentricity in the signals, the localizations can improve significantly. The improvement is expected to be more pronounced for lighter binaries where the number of low-frequency inspiral cycles in the sensitive frequency band is larger. In the absence of a residual eccentricity in the signals, the accuracy of the method falls back to that of the non-eccentric localizations on average.
The basic principle of the localization methods used in this work is to assume that the intrinsic source parameters obtained from matched-filter searches are the true values themselves  [^1] Alternately, this assumption can be lifted toward a full Bayesian parameter estimation, thus increasing the parameter dimensionality and the cost of computing the likelihoods. Several optimized approaches already implement this idea without including eccentricity (Morisaki and Raymond, 2020; You et al., 2021; Dax et al., 2021; Pathak et al., 2024).. In practice, the detection algorithms provide point estimates of these parameters that can carry large uncertainties, e.g. for eccentric sources (Pal and Nayak, 2024). Ideally, the set of true source parameters maximizes the recovered signal-to-noise ratio (SNR), however, as long as the any other combination forming a detection candidate, which nearly maximizes the SNR, can still be used. It has been shown that such a combination can be utilized for the purpose of rapid localization, laying the basis for the BAYESTAR algorithm, which produces skymaps that are almost as accurate as that of a full parameter estimation. This has been demonstrated for sources with small aligned-spins (Singer and Price, 2016).
@@ II Simulated searches
This section briefly describes the processes used to recover simulated signals from simulated data. We create a simulated population of BNS systems with the signal parameters given in Table 1. To generate the signals, we use the TaylorF2Ecc waveform model (Moore et al., 2016). For any given detector, a signal is projected from the random sky location and is deposited in Gaussian noise colored with the simulated PSD for the detector. To recover injections in this work, we broadly make the following two changes to the PSO-based detection framework described in (Pal and Nayak, 2024). These changes are made primarily to simplify simulating the search process and to efficiently tackle a larger search parameter space.
@@
Figure 2: The coincident SNRs recovered by the PSO searches for simulated signals in the HL network having O4-like sensitivities. The searches are carried out in the frequency range of 20-512 Hz using the same datasets containing signals from spin-aligned eccentric BNSs immersed in Gaussian noise. The injected SNRs are computed at the injected parameters with a high-frequency cutoff of 2048 Hz. We use a fixed template-sampling for both the searches as described in the text. The eccentric search recovers all the injections that are also recovered by the aligned-spin search. The commonly found injections are shown here.
@@ III Localization algorithm
Figure 3: PP-plots for the 3D localization parameters shown for the eccentric analysis of the simulated sources described in the Table. 1. These are obtained without any non-trivial correction factor (or with a corresponding default value of 1), which is described later in the text. The gray regions represent the $1$, $2$, and $3\sigma$ uncertainty estimates obtained from beta distribution.
Figure 4: Accuracy (left) and self-consistency (right) of the localization areas compared for the simulated eccentric sources. The solid curves are obtained using the estimated intrinsic source parameters with the PSO-based searches. The dashed curves are obtained using the injected source parameters. The probability plots on the right show that the credible sky areas are consistent at a 99.7% confidence.
Figure 5: Cumulative historgrams showing the distribution of a credible (90%) area for the eccentric injections obtained with the eccentric and non-eccentric analyses.
Figure 6: Skymaps for an eccentric injection obtained with the aligned-spins only (left) and the eccentricity-optimized (right) localization algorithms. The injection is a randomly chosen (1.93, 1.88) $M_{\odot}$ source located at about 65 Mpc, having an eccentricity of 0.1 and nominal aligned-spin values. The localizations are computed following sections Sec. II and III with simulated HLV network. The true location of the simulated source is marked with a star. Notably, we observe that the localization areas at the given credible levels are significantly improved with the eccentricity-optimized algorithm. Further, note the existence of two probability modes (left) compared to a single mode (right) at 90% credible level. Tracking skymaps with multiple separated modes could be resource-demanding process in the EM follow-up communities.
Figure 7: Dependence of the localization improvement on the eccentricity in terms of the 90% confidence area for the eccentric injections obtained with the eccentric and non-eccentric analyses described in the text.
Once a detection has been made, the strain data around the event time from the observing detectors become of particular interest. Here, the localization algorithm receives a sufficient duration of the strain data containing the entire in-band signal determined by its chirp time. The chirp time is calculated using the chirp mass estimate of the source during its detection and the low-frequency cutoff. Note that the PSO search provides a detection candidate corresponding to the template point that optimizes the SNR computed for the detector network. Given the above information, it is possible to assign a probability that the source is located at any given part of the sky. As indicated in (Singer and Price, 2016), the localization uncertainty is only semi-dependent on the uncertainties of the intrinsic source parameters. Thus we conduct a semi-Bayesian analysis by fixing the intrinsic parameters to that obtained as point estimates from the optimized template  [^4] Both the detection and the localization methods reported in this work will need to account for the earth’s rotation effects during a longer inspiral, depending upon the low-frequency cutoff of the analyses. However, note that the output from only a single (optimized) template bridges the two algorithms. Thus, as long as the detection algorithm reports an optimal candidate, a rotation-optimized localization algorithm within the current scheme is also expected to be optimal. We skip such investigations for a future study..
After fixing all the intrinsic parameters, we are left with the extrinsic parameters as shown in the Table. 1. We stochastically draw likelihood samples from the following set of parameters,
@@
Note that the parameters given in Eq. 1 are intended to aid electromagnetic follow-up efforts. Since the injected population has uniformly distributed signal parameters in their respective domains, here we use a uniform prior for each of the parameters. When the injected population is more astrophysically informed, it may be reasonable to use a similar distribution for the prior, see for example (Fairhurst, 2018). This can be achieved with minimal changes in the analysis configuration. Here we choose to marginalize the remaining extrinsic parameters. When the form of the signal can be adequately represented in the dominant (2, 2) mode, the uniform prior over $\varphi$ leads to a simple analytic marginalization of the posterior. The marginalizations over $t_{c}$ and $\psi$ use a uniform grid of $1000$ points in their respective prior spaces and require interpolation of the likelihood followed by explicit numeric marginalization over the resulting array of points. The posterior samples are obtained with the dynesty sampler (Speagle, 2020), configured with nlive $=1000$ and dlogZ $=0.1$. This is accomplished within $25$ seconds using $\sim 32$ CPU cores. A probability-probability (pp)-plot derived from the posterior samples for each of the three-dimensional (3D) localization parameters is shown in Fig. 3. Note that the posterior samples are obtained in the flat $\alpha-\delta$ coordinates. To compare with the localizations from BAYESTAR (Singer and Price, 2016), the samples are rendered into multi-ordered HEALPix format (Fernique et al., 2015; Górski et al., 2005; Zonca et al., 2019). Assigning probabilities to the pixels tiling the sky using the distribution of samples requires an additional postprocessing step (lig, ).
[…]
To assess the self-consistency of the localizations, we use the pp-plot to describe the credibility of the localization areas on the skymaps. A given credible localization patch on the skymap is determined by the group of pixels whose probabilities add up to the desired credible level, starting at the most probable pixel. The pp-plot is calculated by obtaining the fractions of the localized population containing the true location at all credible levels. This is shown in Fig. 4 (right). We note that each configuration of localization described above requires a correction factor on the likelihood computations to obtain a self-consistent pp-plot discussed further in Appendix A. Thus, on average, the localization areas obtained with these methods can be considered reliable for the simulated population. We further note that the precision of the localization area at 90% confidence are improved for the eccentric analysis as compared to the non-eccentric one, obtained for the commonly found injections, as shown in Fig. 5. An example of the localization differences is illustrated in Fig. 6. The general dependence of the improved precision on the eccentricity for the simulated population can be seen in Fig. 7. While the improvement generally scales with eccentricity, it is further dependent on several other factors. The in-band signal, which depends on the frequency cutoffs and the chirp mass, primarily dictates how critical a given value of eccentricity is. The improvement could also be dependent on the unoptimized area itself, which depends on the distance and the orientation of the source as discussed earlier. A discussion on the distance estimates under different priors is presented in Appendix B.
@@
Figure 9: Variation of the loss in the recovered SNR by a simulated aligned-spin search with the injected eccentricity and the total mass of the system in O5-like sensitivity.
Figure 10: Illustration of early-warning skymaps for an eccentric injection. Here the injected eccentricity is 0.05 for a (2.76, 1.43) $M_{\odot}$ source represented with a star. The left and the right panels show the aligned-spin and the eccentricity-informed skymaps; top, middle and bottom panels show the skymaps obtained till 32, 64 and 512 Hz respectively. Given the initial probability densities of the aligned-spin skymaps (the most probable locations are shown with solid pluses), a telescope should begin tracking at the lower left part in the skymap. However, the telescope ideally needs to be repointed as the final skymap indicates to a different location in the sky. The eccentric localizations do not require significant repointings in this case and the most probable locations are relatively more accurate. Here the quoted confidence areas are largely similar in both types of skymaps but we note the existence of such an effect in addition to the cases where the localization areas are themselves numerically improved. A similar effect is reported for non-eccentric sources in (Tohuvavohu et al., 2024).
Figure 11: PP-plots obtained from the localization of commonly found eccentric injections by the EW searches. Plots for only the simulated O4 scenario are shown. Note that the current approach of localization is optimized toward the eccentric injections.
@@
Table 2: Approximate runtimes for generating event candidates from 256 seconds of Gaussian HL data stretch using the spin-aligned eccentric PSO search, assuming that the data exist on the disk. The runtimes for CPU and GPU are obtained on a single AMD EPYC 9554 and NVIDIA Tesla V100 respectively utilizing all available cores. The GPU implementation is currently optimized for early-warning analyses only. The implementation of the early-warning localizations is not optimized and thus, have similar runtimes as for the full-bandwidth localizations.
Figure 12: Difference between the injected $t_{c}$ of the simulated signals and the mean of the estimated arrival times at H1 and L1 obtained from the PSO based EW searches. The number of injections are out of the 10000 injections when the SNR of the triggers crossed 7 at both the detectors. It is highly unlikely that Gaussian noise can surpass this arbitrary threshold in any detector. The plot indicates that a coincident window of $\sim$ 15 ms is insufficient to capture all coincidences in the detector triggers of astrophysical origin.
@@ IV Early-warning scenarios
Figure 13: Accuracy (left) and self-consistency (right) of the localization areas compared for the simulated aligned-spin sources. We note that in the absence of eccentricity in the signals, the localizations accuracy of the current technique is comparable to that of BAYESTAR. Similar to Fig. 4, the localization uncertainties are shown to be fairly self-consistent at 99.7% confidence.
@@
The commonly found injections are then localized in a similar fashion as described in Section III. The accuracies of the localization areas are compared for the observing scenarios as shown in Fig. 8. We observe that the eccentric localization is consistently more accurate than the non-eccentric one in all cases. However, the improvements are generally more prominent in O5-like sensitivities. We primarily attribute this to the greater SNR gain from the numerous low-frequency inspiral cycles while considering eccentricity at the detection stage, as indicated in Fig. 9. We argue that at very low data sampling rates, the SNR consistency across the detectors would carry relatively significant localization information, given the large errors in time and related phase measurements. An interesting comparison of the skymaps at different sampling frequencies is illustrated in Fig. 10. The self-consistency of the sky areas obtained from the EW localizations are validated in Fig. 11.
In the above discussion, we have assumed a 15 ms time-window to test coincidences between triggers from LIGO Hanford (H1) and LIGO Livingston (L1). However, since the strain data for the EW analyses are critically downsampled, the timing uncertainty of the triggers of even astrophysical origin can result into non-coincidence, though they pass an SNR theshold criterion. This is demonstrated for the eccentric injections recovered with SNRs of at least 7 in both detectors as shown in Fig. 12. Note that the improved SNRs can lower the timing uncertainties in general, see for example (Fairhurst, 2011). We further observe that while an eccentric recovery obtains unbiased timing measurements on average, the aligned-spin only searches can report triggers with larger timing errors and/or with bias. Such errors are expected to propagate into the localizations which we do not explore here further. Refer to Appendix C for investigations on the background event rates resulting from increased coincidence window.